# Pay for The Second-Best Service: A Game-Theoretic Approach Against Dishonest LLM Providers

### v3

@@ Abstract.
Mechanism Design, Large Language Models, Online Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the ACM Web Conference 2026; April 13-17, 2026; Dubai, United Arab Emirates11footnotetext: Shanghai Qi Zhi Insititute 2Tsinghua University 3Xiongan AI Institute 4ShanghaiTech University 5Institute of Information Engineering, Chinese Academy of Sciences, Beijing, China 6Beijing Institute for General Artificial Intelligence 7Shanghai University of Finance and Economics
@@ 1. Introduction
In recent years, large language models (LLMs) have demonstrated capabilities across a multitude of domains (thirunavukarasu2023large; hong2024metagpt; balunovic2025matharena). Following the release of ChatGPT (chatgpt2022), there has been a rapid and widespread adoption of LLMs among the general public, embedding them into a variety of commonplace activities. The inference processes of these models are predominantly controlled by companies, which provide access to individual users through interactive interfaces and Application Programming Interfaces (APIs). The growing user base has also triggered the development of a secondary market (e.g., TogetherAI (togetherai2023)), where providers aggregate resources to offer more affordable services.
This market expansion introduces a critical vulnerability: the potential for service provider dishonesty. Since users interact with LLMs in a “black-box” manner, a clear economic incentive exists for providers to substitute the advertised models with less capable alternatives to reduce their own operational costs. Such deceptive practices can manifest in various forms. For instance, when user is asking for a large model like Qwen3-235B-A22B (yang2025qwen3technicalreport), a provider might secretly switch to a smaller one, such as Qwen3-30B-A3B. It might also deploy a quantized version (e.g., using INT8 or FP8) (gao2025model). Additionally, providers might dynamically adjust a model’s reasoning effort (as in reasoning models like OpenAI o1(o12024)) without user consent. The pay-per-token API billing also introduces new risks, allowing providers to inflate charges by inserting superfluous tokens that are difficult for users to detect (kumar2025overthinkslowdownattacksreasoning).
@@
Existing research has approached this challenge from a technical detection standpoint. Efforts include auditing model outputs for semantic behavioral drift (chen2023chatgptsbehaviorchangingtime; 10.1145/3630106.3659047) or token-level consistency (miao-fang-2025-user; sun2025coincountinginvisiblereasoning), and examining the inherent discrepancy across different models (gao2025model; sun2025idiosyncrasieslargelanguagemodels; cai2025gettingpayforauditing). In this work, we tackle the problem through the lens of algorithmic game theory and computational economics. In other words, we propose a mechanism in which the providers would compete with each other, thus protecting user utility.
@@ 1.1. Our Contributions
We formulate a mechanism design problem in which providers determine their strategies based on the user’s delegation mechanism. We characterize the providers’ action space, which comprises cost and token sequence reporting. Our proposed model, which draws upon both mechanism design theory (nisan1999algorithmic) and contract theory (10.1561/0400000113), addresses a novel problem that has not been extensively studied in prior literature.
To better study the relationship between provider actions and outcomes, we also model the performance of LLMs. This specifies how cost control affects the expected reward and output sequence length.
@@
We provide a theoretical guarantee under the continuous cost control model. Inspired by (braverman2019multi) in the online learning domain, we first propose an impossibility theorem, proving that there is no $o(T)$-approximately incentive compatible mechanism that can guarantee a first-best user utility (formally defined later) asymptotically. Based on this, our objective is to propose a mechanism that guarantees a utility close to the second-best user utility.
@@ 2. Related Work
Numerous studies have investigated methods for auditing the disparities between the expected and actual outputs of commercial LLMs from various perspectives. chen2023chatgptsbehaviorchangingtime; 10.1145/3630106.3659047 investigates how LLMs change their behavior over time. sun2025idiosyncrasieslargelanguagemodels explores the idiosyncrasies present in the text generated by different language models, finding that these unique characteristics can persist even after the text has been rewritten, translated, or summarized by external LLMs. sun2025coincountinginvisiblereasoning studies the auditing of invisible reasoning tokens in reasoning LLMs, specifically examining the existence of token count inflation with low-effort token injection under conditions of increased partial embedding exposure and minimal token visibility. cai2025gettingpayforauditing investigates the use of Trusted Execution Environments (TEEs), a hardware architecture, to provide cryptographic guarantees of model integrity. gao2025model propose a concise and effective statistical testing scheme that quantifies changes in the output distribution due to quantization and watermarking by testing for the Maximum Mean Discrepancy between distributions.
@@
The advancement of large language models brings new economic scenarios and problems. 10.1145/3589334.3645511 introduce the token auction model, a mechanism design framework that allows multiple LLM agents to influence generated content through bids. bergemann2025economicslargelanguagemodels provide economic frameworks for token pricing, and 10.1145/3589334.3645366; sun2024mechanismdesignllmfinetuning analyse the game-theoretic structure of fine-tuning and RLHF respectively. saig2024incentivizing investigates how to incentivize LLM service providers to generate high-quality texts via cost-robust contracts.
@@
Algorithms from online learning can be utilized to control LLMs under uncertainty. zhang2023banditrlhf frame RLHF as a contextual bandit and use Thompson sampling to select the most informative comparison queries. chen2024exploration study the exploration–exploitation trade-off when an LLM generates candidate answers with an $\epsilon$-greedy wrapper. wang2025auditingmab treat safety auditing as an adversarial multi-armed bandit. Recent work (bouneffouf2025surveymultiarmedbanditsmeet) systematically explore future directions for combining multi-armed bandits with LLMs. Drawing inspiration from mechanism design in (braverman2019multi), we propose a novel model to address problems within the LLM scenario. This new context, in turn, presents new challenges for theoretical guarantees.
@@ 3.1. The User-Provider Delegation Game
    Cost control: For query $t$, the provider $d_{t}$ can secretly set $c_{t}\in[0,p_{d_{t}}]$ to a value other than the truthful one. For example, the provider could choose a cheaper, lower-quality model or sub-model to generate the response, or resort to quantization (gong2024surveylowbitlargelanguage).
@@
    Token sequence reporting: For query $t$, the provider $d_{t}$ can report a token sequence $\tau^{\prime}_{t}$ that is longer than the actual generated sequence $\tau_{t}$ to inflate its payment. This can be done by adding inconsequential reasoning tokens, or by using techniques like watermarking (zhao2024sokwatermarkingaigeneratedcontent) to add tokens that are difficult for humans to discern. We assume such misreporting has negligible cost. [^2] Note that how such inflation is realized is out of scope for this work.
We assume that, once the user commits to a certain mechanism (made public to everyone), the provider determines its strategy, which remains fixed throughout the execution of the mechanism, regardless of the strategies adopted by other providers. We further assume that, given a history context and a current query, the strategy would choose an action in a deterministic manner.
In reality, the strategy space is discrete and finite (e.g., the provider can only choose from a limited number of models). In our theory, for mathematical convenience, we assume a continuous strategy space, where $c_{t}$ and the reported length $|\tau^{\prime}_{t}|$ are both real numbers.
@@ 3.2. Goals of Mechanism Design
$S_{-i}=(S_{1},\dots,S_{i-1},S_{i+1},\dots,S_{K})$ of other models, we have
$u_{i}^{p}(M,S_{i},S_{-i})\geq u_{i}(M,S_{i}^{\prime},S_{-i}).$
@@
where $\mu_{i}^{r}\in[0,R]$ and $\mu_{i}^{l}\in[1,L]$ is the provider’s underlying expected reward and length when incurring the truthful cost, respectively.
Assume provider $i^{*}$ can deliver $u_{FB}$. Similarly, we define the second-best utility as follows:
@@ 5.1. Notations
To formally link a provider’s actions to outcomes, we model the relationship between cost, output length, and reward.z
@@
We assume providers have an internal belief about their cost-performance trade-offs, modeled by two functions, $g_{i}(c)$ and $h_{i}(c)$, where $c\in[0,p_{i}]$, under the boundary scondition $g_{i}(p_{i})=\mu_{i}^{l}$ and $h_{i}(p_{i})=\mu_{i}^{r}$ (both $\mu_{i}^{l}$ and $\mu_{i}^{r}$ are defined in Section 3). The expected length is defined such that $\mathbb{E}[|\tau_{i,c}|]=g_{i}(c)$. The expected reward is defined such that $\mathbb{E}[v_{i,c}]=h_{i}(c)$. We assume they are Lipschitz continuous.
@@ 5.2. Main Results
[The impossibility of the first-best user utility] For any $\xi\in(0,1)$, there is no $o(T)$-approximate incentive compatible mechanism can guarantee an expected user utility of $\xi u_{FB}+(1-\xi)u_{SB}$, where $u_{FB}$ and $u_{SB}$ are the first-best and the second-best user utility (defined in Section 3.2).
@@
  • $\frac{\mathop{}\!\mathrm{d}h_{i}}{\mathop{}\!\mathrm{d}c}(c)-p_{i}\frac{\mathop{}\!\mathrm{d}g_{i}}{\mathop{}\!\mathrm{d}c}(c)\geq\gamma$,
@@
This theorem relies on two assumptions. First, we assume a positive correlation between a provider’s incurred cost and the utility delivered to the user. This assumption is justifiable within our context, i.e., for each provider, in their own choices, LLMs of larger cost usually yield better performance. The second assumption is that, as cost-per-token rises, the growth in reward to the user must outpace the increase in total cost.
@@
[Second-best user utility guarantee] If all providers action according to Theorem 5.2, the user could get an expected utility of at least $u_{SB}-O(T^{1-\epsilon}\log T+T^{2\epsilon})$.
@@ 6. Experiments
To verify the correctness of our mechanism, we conduct simulation experiments. This section presents the experiments in detail.
@@ 6.1. Setup
Our scenario involves three independent providers, each offering three models. When a user delegates a query to a provider, the provider selects a model according to its strategy, returns the results to the user, and reports the model and the number of tokens used (which may be inaccurate), thereby charging the user. In addition to our proposed optimal strategy, we consider several other strategies, as shown in Table 1. Provider 1 employs three models: gpt-5-high (gpt-5), deepseek-r1 (deepseekai2025deepseekr1incentivizingreasoningcapability), and gpt-5-medium. Provider 2 employs o3-mini (o3mini), o1-mini (o12024), and deepseek-r1. Provider 3 offers claude-4.0 (claude), o1-mini, and deepseek-r1. All model pricing information is obtained from the official websites of the respective models.
@@
We select the maximum flow task from NLGraph (wang2023can) as the dataset for our experiments for the following reasons: (1) the maximum flow task enables the generation of a large number of experimental samples; (2) current LLMs’ performance on the task has not yet reached saturation. Following wang2023can, we use the “partial credit” as the evaluation metric for a single query (which is also the user’s reward), which ranges from 0 to 1, with higher values indicating better performance.
@@
In our main experiment, we set $T$ to 1 million. To reduce costs and accelerate the process, we run 2,000 results for each LLM on NLGraph and repeatedly sample these 2,000 results during the experiment. We set $\epsilon$ to 0.3, and $L$ to 38058 (the maximum output length in our samples) for each provider $i$. To verify the user utility guarantee, we also conduct a supplementary experiment by varying $T$ from 1 million to 2 million.
@@ 7. Limitation
First, our exploration of the model is limited. Although we have established a novel game model, there is considerable room for further investigation. Our theory does not delve into the mechanism design for scenarios where the token sequence length and the action space for cost control are discrete. Furthermore, our proposed mechanism requires the provider to have complete prior knowledge of the model’s capabilities. Further, we do not consider potential collusion, which is also a challenge to mechanism design.
Second, the model itself has the potential for expansion to address more complex real-world situations. For instance, in practice, a provider could maliciously shrink the output token sequence. It is also possible for a provider to increase the length of the token sequence through prompt injection without affecting the LLM’s performance, while still honestly reporting the token sequence.
@@ 8. Conclusion and Future Work
This work confronts the critical issue of service provider dishonesty within the black-box LLM API market, by shifting the focus from technical detection to the lens of algorithmic game theory. We formalize this challenge as a user-provider delegation game, for which we establish a foundational impossibility theorem. In response, we propose an approximately incentive-compatible mechanism that guarantees the user a near-optimal utility, the effectiveness of which is validated through comprehensive simulations grounded in real-world API parameters.
To the best of our knowledge, this work represents a foundational step towards ensuring a more transparent and trustworthy market for large language model services. Future research could extend our model to encompass multiple users, fostering a more competitive and fair market environment. Another valuable problem is to assume the user have a budget constraint, which is a more realistic assumption and a challenge for our model. By establishing a framework grounded in economic principles, we hope to inspire further research at the intersection of artificial intelligence and mechanism design, fostering a healthier and more reliable ecosystem for all users of large language models.

### v4

@@ Abstract.
Mechanism Design; Large Language Models; Online Learning
copyright: acmlicensedjournalyear: 2026copyright: ccconference: the ACM Web Conference 2026; April 13-17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792446ccs: Theory of computation Algorithmic mechanism designccs: Computing methodologies Natural language generationccs: Computing methodologies Online learning settings
@@ 1. Introduction
In recent years, large language models (LLMs) have demonstrated capabilities across a multitude of domains (Thirunavukarasu et al., 2023; Hong et al., 2024; Balunović et al., 2025). Following the release of ChatGPT (OpenAI, 2022), there has been a rapid and widespread adoption of LLMs among the general public, embedding them into a variety of commonplace activities. The inference processes of these models are predominantly controlled by companies, which provide access to individual users through interactive interfaces and Application Programming Interfaces (APIs). The growing user base has also triggered the development of a secondary market (e.g., TogetherAI (TogetherAI, 2023) and OpenRouter (OpenRouter, 2025)), where providers aggregate resources to offer more affordable services.
This market expansion introduces a critical vulnerability: the potential for service provider dishonesty. Since users interact with LLMs in a “black-box” manner, a clear economic incentive exists for providers to substitute the advertised models with less capable alternatives to reduce their own operational costs. Such deceptive practices can manifest in various forms. For instance, when user is asking for a large model like Qwen3-235B-A22B (Yang et al., 2025), a provider might secretly switch to a smaller one, such as Qwen3-30B-A3B. It might also deploy a quantized version (e.g., using INT8 or FP8) (Gao et al., 2025). Additionally, providers might dynamically adjust a model’s reasoning effort (as in reasoning models like OpenAI o1 (OpenAI, 2024a)) without user consent. The pay-per-token API billing also introduces new risks, allowing providers to inflate charges by inserting superfluous tokens that are difficult for users to detect (Kumar et al., 2025).
@@
Existing research has approached this challenge from a technical detection standpoint. Efforts include auditing model outputs for semantic behavioral drift (Chen et al., 2023; Eyuboglu et al., 2024) or token-level consistency (Miao and Fang, 2025; Sun et al., 2025a), and examining the inherent discrepancy across different models (Gao et al., 2025; Sun et al., 2025b; Cai et al., 2025). In this work, we tackle the problem through the lens of algorithmic game theory and computational economics. In other words, we propose a mechanism in which the providers would compete with each other, thus protecting user utility.
@@ 1.1. Our Contributions
We formulate a mechanism design problem in which providers determine their strategies based on the user’s delegation mechanism. We characterize the providers’ action space, which comprises cost and token sequence reporting. Our proposed model, which draws upon both mechanism design theory (Nisan and Ronen, 1999) and contract theory (Dütting et al., 2025), addresses a novel problem that has not been extensively studied in prior literature. To better study the relationship between provider actions and outcomes, we also model the performance of LLMs. This specifies how cost control affects the expected reward and output sequence length.
@@
We provide a theoretical guarantee under the continuous cost control model. Inspired by (Braverman et al., 2019) in the online learning domain, we first propose an impossibility theorem, proving that there is no $o(T)$-approximately incentive compatible mechanism that can guarantee a first-best user utility (formally defined later) asymptotically. Based on this, our objective is to propose a mechanism that guarantees a utility close to the second-best user utility.
@@ 2. Related Work
Numerous studies have investigated methods for auditing the disparities between the expected and actual outputs of commercial LLMs from various perspectives. Chen et al. (2023); Eyuboglu et al. (2024) investigates how LLMs change their behavior over time. Sun et al. (2025b) explores the idiosyncrasies present in the text generated by different language models. Sun et al. (2025a) studies the auditing of invisible reasoning tokens in reasoning LLMs, specifically examining the existence of token count inflation with low-effort token injection. Cai et al. (2025) investigates the use of Trusted Execution Environments (TEEs), a hardware architecture, to provide cryptographic guarantees of model integrity. Gao et al. (2025) propose a concise and effective statistical testing scheme that quantifies changes in the output distribution due to quantization and watermarking by testing for the Maximum Mean Discrepancy between distributions. Recent works (Velasco et al., 2025b, a) systematically explore how LLM providers misreport token usage to gain additional utility under identical string representations, as well as how to audit such behaviors.
@@
The advancement of large language models brings new economic scenarios and problems. Dütting et al. (2024) introduce the token auction model, a mechanism design framework that allows multiple LLM agents to influence generated content through bids. Bergemann et al. (2025) provide economic frameworks for token pricing, and Laufer et al. (2024); Sun et al. (2024) analyse the game-theoretic structure of fine-tuning and RLHF respectively. Saig et al. (2024) investigates how to incentivize LLM service providers to generate high-quality texts via cost-robust contracts.
@@
Algorithms from online learning can be utilized to control LLMs under uncertainty. Zhang et al. (2023) frame RLHF as a contextual bandit and use Thompson sampling to select the most informative comparison queries. Chen et al. (2024) study the exploration–exploitation trade-off when an LLM generates candidate answers with an $\epsilon$-greedy wrapper. Wang et al. (2025) treat safety auditing as an adversarial multi-armed bandit. Recent work (Bouneffouf and Feraud, 2025) systematically explore future directions for combining multi-armed bandits with LLMs. Drawing inspiration from mechanism design in (Braverman et al., 2019), we propose a novel model to address problems within the LLM scenario. This new context, in turn, presents new challenges for theoretical guarantees.
@@ 3.1. The User-Provider Delegation Game
    Cost control: For query $t$, the provider $d_{t}$ can secretly set $c_{t}\in[0,p_{d_{t}}]$ to a value other than the truthful one. For example, the provider could choose a cheaper, lower-quality model or sub-model to generate the response, or resort to quantization (Gong et al., 2024).
@@
    Token sequence reporting: For query $t$, the provider $d_{t}$ can report a token sequence $\tau^{\prime}_{t}$ that is longer than the actual generated sequence $\tau_{t}$ to inflate its payment. This can be done by adding inconsequential reasoning tokens, or by using techniques like watermarking (Zhao et al., 2024) to add tokens that are difficult for humans to discern. We assume such misreporting has negligible cost. [^2] Note that how such inflation is realized is out of scope for this work.
In reality, the strategy space is discrete and finite (e.g., the provider can only choose from a limited number of models). In our theory, for mathematical convenience, we assume a continuous strategy space [^3] In the discrete setting, our mechanism would introduce an additional approximation error of a single token cost (around $10^{-6}$ USD in practice) per query. It leads to a cumulative error of $O(T)$ over $T$ queries. Since the error per query is negligible, we adopt the continuous version in our model., where $c_{t}$ and the reported length $|\tau^{\prime}_{t}|$ are both real numbers.
@@ 3.2. Goals of Mechanism Design
$S_{-i}=(S_{1},\dots,S_{i-1},S_{i+1},\dots,S_{K})$ of other providers, we have
$u_{i}^{p}(M,S_{i},S_{-i})\geq u_{i}^{p}(M,S_{i}^{\prime},S_{-i}).$
@@
where $\mu_{i}^{r}\in[0,R]$ and $\mu_{i}^{l}\in[1,L]$ is the provider’s underlying expected reward and length of token sequence when incurring the truthful cost, respectively.
Assume provider $i^{*}$ can deliver $u_{FB}$. Similarly, we define the second-best user utility as follows:
@@ 5.1. Notations
To formally link a provider’s actions to outcomes, we model the relationship between cost, output length, and reward.
@@
We assume providers have an internal belief about their cost-performance trade-offs, modeled by two functions, $g_{i}(c)$ and $h_{i}(c)$, where $c\in[0,p_{i}]$, under the boundary condition $g_{i}(p_{i})=\mu_{i}^{l}$ and $h_{i}(p_{i})=\mu_{i}^{r}$ (both $\mu_{i}^{l}$ and $\mu_{i}^{r}$ are defined in Section 3). The expected length is defined such that $\mathbb{E}[|\tau_{i,c}|]=g_{i}(c)$. The expected reward is defined such that $\mathbb{E}[v_{i,c}]=h_{i}(c)$. We assume they are Lipschitz continuous.
@@ 5.2. Main Results
[The impossibility of the first-best user utility] For any $\xi\in(0,1)$, there is no $o(T)$-approximately incentive compatible mechanism can guarantee an expected user utility of $\xi u_{FB}+(1-\xi)u_{SB}$, where $u_{FB}$ and $u_{SB}$ are the first-best and the second-best user utility (defined in Section 3.2).
@@
  • $\frac{\mathop{}\!\mathrm{d}h_{i}}{\mathop{}\!\mathrm{d}c}(c)-p_{i}\frac{\mathop{}\!\mathrm{d}g_{i}}{\mathop{}\!\mathrm{d}c}(c)\geq\gamma$ for some $\gamma>0$,
@@
This theorem relies on two assumptions. First, we assume a positive correlation between a provider’s incurred cost and the utility delivered to the user. This assumption is justifiable within our context, i.e., for each provider, in their own choices, LLMs of higher cost always yield better performance. The second assumption is that, as cost-per-token rises, the growth in reward to the user must outpace the increase in total cost.
@@
[Second-best user utility guarantee] If all providers act according to Theorem 5.2, the user could get an expected utility of at least $u_{SB}-O(T^{1-\epsilon}\log T+T^{2\epsilon})$.
@@ 6. Experiments
To demonstrate the efficacy of our mechanism, we conduct simulation experiments. This section presents the experiments in detail.
@@ 6.1. Setup
Our scenario involves three independent providers, each employing three models. We use the price of the best model as the provider’s price, and the price of each model as the provider’s different cost options. When a user delegates a query to a provider, the provider selects a model according to its strategy, returns the output token sequence to the user, and reports the number of tokens used (which may be inaccurate), thereby charging the user. In addition to our proposed strategy, we consider several other strategies, as shown in Table 1. Provider 1 employs three models: gpt-5-high (openai, 2025), deepseek-r1 (DeepSeek-AI et al., 2025), and gpt-5-medium. Provider 2 employs o3-mini (OpenAI, 2024b), o1-mini (OpenAI, 2024a), and deepseek-r1. Provider 3 offers claude-4.0 (anthropic, 2025), o1-mini, and deepseek-r1. All model pricing information is obtained from the official websites of the respective models.
@@
We select the maximum flow task from NLGraph (Wang et al., 2023) as the dataset for our experiments for the following reasons: (1) the maximum flow task enables the generation of a large number of experimental samples; (2) current LLMs’ performance on the task has not yet reached saturation. Following Wang et al. (2023), we use the “partial credit” as the evaluation metric for a single query (which is also the user’s reward), ranging from 0 to 1, with higher values indicating better performance.
@@
In our main experiment, we set $T$ to 1 million. To reduce costs and accelerate the process, we run 2,000 results for each LLM on NLGraph and repeatedly sample these 2,000 results during the experiment. We set $\epsilon$ to 0.3, and $L$ to 38,058 (the maximum output length in our samples). To verify the user utility guarantee, we also conduct a supplementary experiment by varying $T$ from 1 million to 2 million.
@@ 7. Limitation
First, our exploration of the model is limited. Although we have established a novel game model, there is considerable room for further investigation. Our theory does not delve into the mechanism design for scenarios where the token sequence length and the action space for cost control are discrete. Furthermore, our proposed mechanism requires all providers to have complete prior knowledge of the model’s capabilities.
Second, our model simplifies several real-world complexities. We assume no collusion between providers; however, drawing a parallel to auction theory, the first-best and second-best providers may have strong incentives to engage in collusive behavior if mutual recognition exists. Additionally, we assume that the user (mechanism designer) acts truthfully within the mechanism; relaxing this assumption could undermine the mechanism. For example, a user might be incentivized to strategically misreport $B$, the number of queries per provider in the exploration phase.
@@ 8. Conclusion and Future Work
This work confronts the critical issue of service provider dishonesty within the black-box LLM API market, by shifting the focus from technical detection to the lens of algorithmic game theory. We formalize this challenge as a user-provider delegation game, for which we establish a foundational impossibility theorem. In response, we propose an approximately incentive-compatible mechanism that guarantees the user a near-optimal utility, the effectiveness of which is validated through comprehensive simulations grounded in real-world API parameters. Our mechanism can be effectively implemented in aggregated API platforms and enterprise clients, both of which have the leverage to contract with different managed service providers.
To the best of our knowledge, this work is the first to propose a game-theoretic framework towards ensuring a more transparent and trustworthy market for large language model services. Future research could extend our model to encompass multiple users, fostering a more competitive and fair market environment. Another valuable problem is to assume the user has a budget constraint, which is a more realistic assumption and a challenge for our model. By establishing a framework grounded in economic principles, we hope to inspire further research at the intersection of artificial intelligence and mechanism design, fostering a healthier and more reliable ecosystem for all users of large language models.